OpenStudy (anonymous):

For the following quadratic functions, find the value(s) of x for the given value of y: a) y=xˆ2+6x+10 when y=1 could you explain how to do this? thanks

myininaya (myininaya):

So we have 1=x^2+6x+10

myininaya (myininaya):

Now to solve quadratic equations you must have one side equal to 0.

myininaya (myininaya):

Do you know how to get one of the sides of the equal sign to be zero?

OpenStudy (anonymous):

well could we minus the 1 from the left side of the equation?

myininaya (myininaya):

Yep. Remember whatever you do to one side, you must do to the other.

myininaya (myininaya):

So what is the resulting equation after subtracting one from both sides?

OpenStudy (anonymous):

0=xˆ2+6x+9

myininaya (myininaya):

Perfect. Do you know how to factor?

OpenStudy (anonymous):

no

myininaya (myininaya):

What about using the quadratic formula?

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

wait with factor is it factorizing? would it give me 0=(x+3)(x+3) ? if not then i don't know about using the quadratic formula either.

myininaya (myininaya):

You factored correctly! :) Now set each of those factors =0 like so x+3=0 or x+3=0 This is the same equation They will have the same solution

OpenStudy (anonymous):

oh so it gives us -3?

myininaya (myininaya):

Yep :)

OpenStudy (anonymous):

wow okay thank you! (: